OpenStudy (anonymous):

the ratio of zebras to monkeys in a circus is 3:5. if there are 56 zebras and monkeys in the circus altogether, then how many zebras are there

OpenStudy (anonymous):

A ratio of 3:5 can be expressed 3x:5x. The sum total of zebras and monkeys is 56, so 3x + 5x = 56 8x = 56 x = 7 So there are 3x = 3(7) = 21 zebras, and there are 5x = 5(7) = 35 monkeys.